A bullet weighing 0.08 lb is fired with a horizontal velocity of 1800 ft/s into the lower end of a slender 19-lb bar of length L 30 in. The bar is initially at rest and h 12 in. Vo References eBook & Resources Section Break Difficulty: Easy 10.00 points Determine the impulsive reaction at C, assuming that the bullet becomes embedded in 0.001 s. (Round the final answer to two decimal places.) The impulsive reaction at C is lb-
Answer #1

Oostb. Bolt/, 1916, h 12n, 1.som 0.001 for Bar 1.30m 2.5 ft. m. 19 = 0.59 16 3/2 393 I 1/1 1/2 (0.5) (25) - 0. 301 29 165.ft

V = 45W - 16 (19-28) = 28.92 Als Vq- (0 asw) = 0.25. ( 19.28) - 4.82 f/5 Honzantal component my to (st) - M - mug ast) = mo (